OpenStudy (anonymous):

Solve the system of equations 5x+3y=58 5x-3y=22

OpenStudy (lgbasallote):

add the simultaneous equations..that will give you 10x = 80 therefore x = 8 just sub in 8 into x in any of the e

OpenStudy (lgbasallote):

*just sun in 8 into x in any of the equations to find y

OpenStudy (lgbasallote):

*just sub*

OpenStudy (s3a):

Substitution method: 5x + 3y = 58 Divide both sides by 5, x + 3/5 y = 58/5 Isolate x, x = 58/5 - 3/5 y shove in x into the x of the second equation, 5x - 3y = 22 5(58/5 - 3/5 y) - 3y = 22 58 - 3y - 3y = 22 58 - 6y = 22 6y = 58 - 22 6y = 36 y = 6